# Accuracy-Robustness Trade Off via Spiking Neural Network Gradient Sparsity Trail

### v2

@@ 4.2 Natural adversarial robustness of SNN
Input discretization (Sharmin et al., 2020) 15.50 6.30 64.40
Hire-SNN (Kundu et al., 2021) 22.00 7.50 65.10
Vanilla VGG-11 (Ding et al., 2022) 5.30 0.02 73.33
SNN-RAT VGG-11 (Ding et al., 2022) 25.86 10.38 70.89
SNN-RAT WRN-16 (Ding et al., 2022) 28.08 11.31 69.32
SEW-ResNet18 (Fang et al., 2021a) 20.81 13.47 42.19
SEW-ResNet34 (Fang et al., 2021a) 19.91 13.25 41.81
SEW-ResNet50 (Fang et al., 2021a) 21.27 16.03 38.91
Spiking-ResNet18 (Hu et al., 2021) 13.36 11.15 23.23
Spiking-ResNet34 (Hu et al., 2021) 2.43 1.64 21.20
Spiking-ResNet50 (Hu et al., 2021) 3.32 2.44 20.56
We benchmark our selected models against several SOTA methods aimed at enhancing the adversarial robustness of spiking neural networks. Specifically, we compare our undefended SNN variants with input discretization (Sharmin et al., 2020), Hire-SNN (Kundu et al., 2021), and SNN-RAT (Ding et al., 2022), all of which have demonstrated notable improvements under adversarial settings. The evaluations will focus on standard attack methods such as FGSM and PGD and all models will be benchmarked using fixed settings for adversarial attacks. Specifically, $\epsilon=8/255$ will be used for both FGSM and PGD, with PGD further configured with $\alpha=0.01$ and $t=7$, consistent with previous researches (Ding et al., 2022).
@@
CIFAR-10 SNN-RAT VGG-11 (Ding et al., 2022) 45.23 21.16 90.74
SNN-RAT WRN-16 (Ding et al., 2022) 50.78 22.71 92.69
SEW-ResNet18 (Fang et al., 2021a) 46.99 34.33 73.78
SEW-ResNet34 (Fang et al., 2021a) 49.20 36.98 74.16
SEW-ResNet50 (Fang et al., 2021a) 51.09 40.55 73.34
CIFAR-100 SNN-RAT VGG-11 (Ding et al., 2022) 25.86 10.38 70.89
SNN-RAT WRN-16 (Ding et al., 2022) 28.08 11.31 69.32
SEW-ResNet18 (Fang et al., 2021a) 20.81 13.47 42.19
SEW-ResNet34 (Fang et al., 2021a) 19.91 13.25 41.81
SEW-ResNet50 (Fang et al., 2021a) 21.27 16.03 38.91
@@
As presented in Table 1, all SEW-ResNet variants exhibit SOTA performance against PGD attacks, achieving up to 16.03% accuracy—an improvement of 4.72% over the previous best method (SNN-RAT WRN-16). For FGSM attacks, SEW-ResNet50 yields a competitive 21.27% accuracy, which is only 4.59% below SNN-RAT and nearly equivalent to the performance of Hire-SNN, leading us to suspect that the benchmarking result was affected by obfuscated gradientAthalye et al. (2018a). In contrast, Spiking-ResNet variants without any defense mechanism do not exhibit similar resilience as deeper residual blocks being constructed, therefore we would only continue benchmarking with SEW-ResNet variants. From Table 2, we can see that this behavior not only consistent on multiple dataset, SEW-ResNet50 was able to outperform all mentioned adversarial defense mechanism with a 0.31% gap on FGSM and 17.84% gap on PGD (compare with SNN-RAT WRN-16).
4.3 Explanation and analysis
@@
As demonstrated in the benchmarking results presented in Section 4.2, we observe that undefended SEW-ResNet exhibit a notable degree of adversarial robustness, achieving performance comparable to SOTA defense methods but Spiking-ResNet does not exhibit the same behavior. This raises a fundamental question: Why do SNNs demonstrate such inherent robustness in the absence of explicit defense mechanisms?
@@
We hypothesize that this phenomenon arises from the inherent sparsity of input gradients in SEW-ResNet , which in turn reduces the efficacy of adversarial perturbations. Supporting this hypothesis, prior work (Liu et al., 2024) has shown that SNN show strong resistance against random perturbation at significant scale, where the distance between random vulnerability and adversarial vulnerability can be upper bounded by the sparsity of the input gradient as:
Theorem 1 (Liu et al.).
@@
Suppose $f$ is a differentiable SNN by surrogate gradients, and $\epsilon$ is the magnitude of an attack, assumed to be small enough. Given an input image $\mathbf{x}$ with corresponding label $y$, the ratio of $\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ and $\rho_{\text{rand}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ is upper bounded by the sparsity of $\nabla_{\mathbf{x}}f_{y}$:
$3\leq\frac{\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})}{\rho_{\text{rand}}(f,\mathbf{x},\epsilon,\ell_{\infty})}\leq 3\|\nabla_{\mathbf{x}}f_{y}(\mathbf{x})\|_{0}.$
As observed from the theorem, adversarial vulnerability $\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ scale positively with non-zero entries in input gradient, we can form a corollary as:
Corollary 1.
@@
The adversarial vulnerability $\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ scale positively with non-zero entries of input gradient $\|\nabla_{\mathbf{x}}f_{y}(\mathbf{x})\|_{0}$.
@@
To validate this explanation in light of the theorem, we conduct an empirical analysis of the input gradient patterns for SEW-ResNet models under adversarial perturbations. Specifically, we compute the average percentage of non-zero entries in the input gradients (as shown in Figure 1). Additionally, we measure the gradient norm to ensure that the sparsity is not a byproduct of shattered gradients, following the recommendations from Athalye et al. (2018a).
@@
$\frac{\partial\text{Conv}(x,\mathbf{w})}{\partial x}$ $=\text{Conv}^{\intercal}\left(\frac{\partial\mathcal{L}}{\partial y},\mathbf{w}\right),$ (5)
$\frac{\partial\text{BN}(y_{i,c})}{\partial x_{i,c}}$ $=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\varepsilon}}\left[\delta_{i,c}-\sum_{j}^{m}\frac{\delta_{j,c}}{m}-\frac{\hat{x}_{i,c}}{m}\sum_{j}^{m}\delta_{j,c}\cdot\hat{x}_{j,c}\right],$
@@
$\frac{\partial\mathcal{H}}{\partial x}$ $=\frac{\alpha}{2\left(1+\left(\frac{\pi}{2}\alpha x\right)^{2}\right)}\quad\text{where $\alpha=2$}$
@@
To quantify these sources, we analyze the gradient expressions of each component within the spike-based element-wise residual blocks. In Equation 5, Conv and $\text{Conv}^{\intercal}$ refer to the forward and transposed convolution operations respectively. $\mathcal{L}$ denotes the loss function, $\mathbf{w}$ the kernel weights, and $y$ the output of the operation. For the batch normalization (BN) gradient, $\gamma_{c}$ and $\beta_{c}$ represent the learned scale and shift parameters for channel $c$, while $\mu_{c}$ and $\sigma_{c}^{2}$ are the mean and variance computed across the batch of size $m$. The notation $\hat{x}_{i,c}$ is the normalized input, and $\delta_{i,c}$ represents the partial derivative of the loss with respect to the BN output. The max pooling gradient selectively propagates gradients only through the maximum element within each pooling region $\mathcal{R}_{p,q}$, and $\mathcal{H}$ denotes the arctangent surrogate gradient of IF neurons, controlled by the slope parameter $\alpha$.
Figure 2: Architecture of SEW residual block proposed by Fang et al. (2021a)
As seen in Equation 5, convolution, batch normalization, and surrogate gradient components are all dependent on input and intermediate values. However, max pooling introduces a structural bottleneck where gradients are non-zero only at the maximum activation locations, and zero elsewhere. These null gradients can propagate through backward pass, create sparsity in earlier layers and is denoted as architectural sparsity. This effect is amplified in SEW-ResNet due to the high coupling nature of residual blocks (illustrated in Figure 2). Thus, replacing max pooling with a less sparse alternative could enhance gradient density. To evaluate this hypothesis, we replaced max pooling operations with average pooling, whose gradient is given by:
$\frac{\partial\text{AvgPool}}{\partial x_{i,j}}=\sum_{(p,q)\;:\;(i,j)\in\mathcal{R}_{p,q}}\;\frac{1}{|\mathcal{R}_{p,q}|}\cdot\frac{\partial\mathcal{L}}{\partial y_{p,q}}$ (6)
@@
Table 3: Comparison of CIFAR-100 and CIFAR-10 validation accuracy of undefended SEW-ResNet variants (Fang et al., 2021a) with average pooling. Best performance per attack on each dataset are denoted in bold. We employed similar attack settings as Table 1, accuracy reduction with respect to Table 2 is denoted in red and accuracy increment is denoted in green. There is a clear trade off in clean accuracy and adversarial robustness.
@@
As shown in Figure 3, input gradient sparsity is significantly reduced when average pooling is used in place of max pooling across all layers. Specifically, sparsity in the pooling layer is completely eliminated, leading to a 37.88% reduction compared to max pooling, while a spiking neuron layers exhibit sparsity as low as 16.52%. This observation is consistent with prior studies. For instance, Fang et al. (2021a) highlighted the importance of gradient magnitude in optimizing SNNs to enhance performance, while Luu et al. (2024b) addressed training challenges by introducing signal concatenation, effectively scaling the gradient norm to improve generalization. Complete elimination of sparsity is also theoretically unachievable due to the inherent nature of SNNs. This residual sparsity arises from the binary nature of the spiking function—typically modeled using the Heaviside function as in Equation 4—and is thus referred to as natural sparsity.
5 How does shared sparsity trail in gradient computation affect performance?
@@
Previous studies have demonstrated a strong relationship between the magnitude of weight gradients and model performance (Luu et al., 2024b; Fang et al., 2021a; Shalev-Shwartz et al., 2017). Building on this foundation, we aim to investigate how sparsity in the input gradients may suppress weight gradients during optimization, thereby influencing overall performance. To prove this statement, we consider the following theorem:
Theorem 2.
@@
As obtained from the derivation results, input gradient $\nabla_{s[t]}p_{\mathbf{w}}(s)$ shared a gradient chain $\frac{\partial p_{\mathbf{w}}}{\partial y_{\mathbf{w}_{2}}}\cdot\frac{\partial y_{\mathbf{w}_{2}}}{\partial z_{\mathbf{w}_{2}}}\cdot\frac{\partial z_{\mathbf{w}_{2}}}{\partial y_{\mathbf{w}_{1}}}\cdot\frac{\partial y_{\mathbf{w}_{1}}}{\partial z_{\mathbf{w}_{1}}}$ with the Jacobian $\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)$ of $\mathbf{w}_{1}$ during backward computation and a shorter chain of $\frac{\partial p_{\mathbf{w}}}{\partial y_{\mathbf{w}_{2}}}\cdot\frac{\partial y_{\mathbf{w}_{2}}}{\partial z_{\mathbf{w}_{2}}}$ with respect to $\mathbf{w}_{2}$, leading us to several observations:
  • Parameters that are placed closer to the initial input signal will shared a longer gradient chain, making sparsity in input gradient have a stronger correlation with parameter gradient [^1] A more illustrative demonstration can be found at Figure 4
We are confident that any sparsity in the shared gradient path propagates backward to both the weight and input gradients. Still, opposite of sparsity, how can we formally quantify the relationship between the density of the weight gradient (specifically of $\mathbf{w}_{1}$, as it is being the closest parameter to input and is highly affected by the shared gradient trail) with respect to input gradient, particularly in the context of Theorem 2? To address this, we establish the following theorem, which provides a precise characterization of their relationship:
Theorem 3.
Given the same condition as Theorem 2, the parameter matrix $\mathbf{w}_{l}\in\mathbb{R}^{d_{1}\times d_{2}}$ at layer $l$ is said to have a row/column density degree of $k^{(i)}_{\mathbf{w}_{l}}$$k^{(j)}_{\mathbf{w}_{l}}$ (or $k^{(i)}_{\mathbf{w}_{l}}$$k^{(j)}_{\mathbf{w}_{l}}-$dense row/column for short) if each of its row/column vector $\mathbf{w}_{l}^{(i)}/\mathbf{w}_{l}^{(j)}$ (where $i\in d_{1}$, $j\in d_{2}$) has at most $k$ non-zero entries $\mathbf{w}^{(ij)}_{l}$, defined as:
@@
then $L_{0}$ norm of weight gradient of $\mathbf{w}_{1}$ can bounded as:
@@
$\mathbb{E}\left[\|\nabla_{s[t]}p_{\mathbf{w}}(s)\|_{0}\right]\leq k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{\mathbf{w}_{1}}$ (16)
@@
$Pr\left(\mathbf{w}^{(in)}_{2}\cdot S^{(nj)}\neq 0\right)\leq\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{d^{2}}\quad\Leftrightarrow\quad\mathbb{E}\left[\|\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)\|_{0}\right]\leq T\cdot\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{d}$ (17)
Divide obtained inequalities from Equation 16 and 17:
$\frac{\mathbb{E}\left[\|\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)\|_{0}\right]}{\mathbb{E}\left[\|\nabla_{s[t]}p_{\mathbf{w}}(s)\|_{0}\right]}\leq\frac{T\cdot\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{d}}{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{\mathbf{w}_{1}}}$ (18)
$\Leftrightarrow$ $\mathbb{E}\left[\|\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)\|_{0}\right]\leq\frac{T}{d}\cdot\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{\mathbf{w}_{1}}}\cdot\mathbb{E}\left[\|\nabla_{s[t]}p_{\mathbf{w}}(s)\|_{0}\right].$

### v3

@@ 4.2 Natural adversarial robustness of SNN over temporal-coded backpropagation
Input discretization (sharmin2020inherent) 15.50 6.30 64.40
Hire-SNN (kundu2021hire) 22.00 7.50 65.10
Vanilla VGG-11 (ding2022snn) 5.30 0.02 73.33
SNN-RAT VGG-11 (ding2022snn) 25.86 10.38 70.89
SNN-RAT WRN-16 (ding2022snn) 28.08 11.31 69.32
SEW-ResNet18 (fang2021deep) 20.81 13.47 42.19
SEW-ResNet34 (fang2021deep) 19.91 13.25 41.81
SEW-ResNet50 (fang2021deep) 21.27 16.03 38.91
Spiking-ResNet18 (hu2021spiking) 13.36 11.15 23.23
Spiking-ResNet34 (hu2021spiking) 2.43 1.64 21.20
Spiking-ResNet50 (hu2021spiking) 3.32 2.44 20.56
We benchmark our selected models on CIFAR10, CIFAR100 (alex2009learning) and CIFAR10-DVS (li2017cifar10) against several SOTA methods aimed at enhancing the adversarial robustness of spiking neural networks. Specifically, we compare our undefended SNN variants with input discretization (sharmin2020inherent), Hire-SNN (kundu2021hire), and SNN-RAT (ding2022snn), all of which have demonstrated notable improvements under adversarial settings. The evaluations will focus on standard attack methods such as FGSM and PGD and all models will be benchmarked using fixed settings for adversarial attacks. Specifically, $\epsilon=8/255$ will be used for both FGSM and PGD, with PGD further configured with $\alpha=0.01$ and $t=7$, consistent with previous researches (ding2022snn). A visualization of the perturbated images are presented in Figure 1.
Figure 1: Example illustration of perturbated images from CIFAR-10 with FGSM and PGD over a network.
@@
CIFAR10 SNN-RAT VGG-11 (ding2022snn) 45.23 21.16 90.74
SNN-RAT WRN-16 (ding2022snn) 50.78 22.71 92.69
SEW-ResNet18 (fang2021deep) 46.99 34.33 73.78
SEW-ResNet34 (fang2021deep) 49.20 36.98 74.16
SEW-ResNet50 (fang2021deep) 51.09 40.55 73.34
CIFAR100 SNN-RAT VGG-11 (ding2022snn) 25.86 10.38 70.89
SNN-RAT WRN-16 (ding2022snn) 28.08 11.31 69.32
SEW-ResNet18 (fang2021deep) 20.81 13.47 42.19
SEW-ResNet34 (fang2021deep) 19.91 13.25 41.81
SEW-ResNet50 (fang2021deep) 21.27 16.03 38.91
CIFAR10-DVS SNN-RAT VGG-11 (ding2022snn) 36.22 19.02 71.77
SNN-RAT WRN-16 (ding2022snn) 40.16 17.11 76.12
SEW-ResNet18 (fang2021deep) 47.19 21.47 75.18
SEW-ResNet34 (fang2021deep) 49.03 18.65 75.29
SEW-ResNet50 (fang2021deep) 48.52 22.34 76.53
@@
As presented in Table 1, all SEW-ResNet variants exhibit SOTA performance against PGD attacks, achieving up to 16.03% accuracy—an improvement of 4.72% over the previous best method (SNN-RAT WRN-16). For FGSM attacks, SEW-ResNet50 yields a competitive 21.27% accuracy, which is only 4.59% below SNN-RAT and nearly equivalent to the performance of Hire-SNN, leading us to suspect that the benchmarking result was affected by obfuscated gradient (athalye2018obfuscated). In contrast, Spiking-ResNet variants without any defense mechanism do not exhibit similar resilience as deeper residual blocks being constructed, therefore we would only continue benchmarking with SEW-ResNet variants. From Table 2, we can see that this behavior not only consistent on multiple dataset, SEW-ResNet was able to outperform almost all mentioned adversarial defense mechanism:
  • SEW-ResNet50 with a 0.31% gap on FGSM and 17.84% gap on PGD (compare with SNN-RAT WRN-16 on CIFAR10),
  • SEW-ResNet50 with a 4.72% gap on PGD (compare with SNN-RAT WRN-16 on CIFAR100),
  • SEW-ResNet34 with a 8.87% gap on FGSM (compare with SNN-RAT WRN-16 on CIFAR10-DVS) and SEW-ResNet50 with a 3.32% gap on PGD (compare with SNN-RAT VGG-11 on CIFAR10-DVS).
Figure 2: Effect of adversarial attacks over multiple perturbation magnitude $\epsilon$ between defended WideResNet16 and vanilla SEW-ResNet18.
We further investigate the adversarial robustness of the benchmarked models across varying perturbation magnitudes, specifically at $\epsilon\in\{2,4,6,8\}/255$. As illustrated in Figure 2, the accuracy of the regularized WideResNet-16 progressively decreases as the input perturbation increases, with performance reductions of approximately $35\%$ under FGSM and $70\%$ under PGD attacks. In contrast, SEW-ResNet exhibits relatively higher vulnerability to smaller perturbations, showing performance degradations of about $25\%$ under FGSM and $35\%$ under PGD, which is consistent with behavior validated by prior work regarding the effect of small scale perturbation on SNN (liu2024enhancing).
4.3 Gradient analysis and theoretical explanation
4.3.1 Gradient analysis
@@
As demonstrated in the benchmarking results presented in Section 4.2, we observe that undefended SEW-ResNet exhibit a notable degree of adversarial robustness, achieving performance comparable to SOTA defense methods but Spiking-ResNet does not exhibit the same behavior. This raises some fundamental questions:
  • Why do SNNs demonstrate such inherent robustness in the absence of explicit defense mechanisms?
@@
We hypothesize that this phenomenon arises from the inherent sparsity of input gradients in SEW-ResNet, which in turn reduces the efficacy of adversarial perturbations while also make small scale perturbation to yield stronger effect. Supporting this hypothesis, prior work has shown that the strength of adversarial attacks can be upper bounded by the sparsity of the input gradient. Specifically, liu2024enhancing established the following theoretical bound:
Theorem 4.1 (liu2024enhancing).
@@
Suppose $f$ is a differentiable SNN by surrogate gradients, and $\epsilon$ is the magnitude of an attack, assumed to be small enough. Given an input image $\mathbf{x}$ with corresponding label $y$, the ratio of adversarial vulnerability $\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ and random vulnerability $\rho_{\text{rand}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ is upper bounded by the density of input gradient $\nabla_{\mathbf{x}}f_{y}$ as:
$3\leq\frac{\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})}{\rho_{\text{rand}}(f,\mathbf{x},\epsilon,\ell_{\infty})}\leq 3\|\nabla_{\mathbf{x}}f_{y}(\mathbf{x})\|_{0}.$ (5)
@@
where $\rho_{\text{adv}}(f,\mathbf{x},\epsilon,\ell_{\infty})$ scale linearly with non-zeros entries ($L_{0}$ norm) of input gradient $\|\nabla_{\mathbf{x}}f_{y}(\mathbf{x})\|_{0}$.
Proof for this theorem is denoted in Appendix A.
@@
To validate this explanation in light of the theorem, we conduct an empirical analysis of the input gradient patterns for SEW-ResNet models under adversarial perturbations. Specifically, we compute the average percentage of non-zero entries in the input gradients (as shown in Figure 3). Additionally, we measure the gradient norm to ensure that the sparsity is not a byproduct of shattered gradients, following the recommendations from athalye2018obfuscated.
@@
$\frac{\partial\textbf{Conv}(x,\mathbf{w})}{\partial x}$ $=\textbf{Conv}^{\intercal}\left(\frac{\partial\mathcal{L}}{\partial y},\mathbf{w}\right),$ (6)
$\frac{\partial\textbf{BN}(y_{i,c})}{\partial x_{i,c}}$ $=\frac{\gamma_{c}}{\sqrt{\sigma_{c}^{2}+\varepsilon}}\left[\delta_{i,c}-\sum_{j}^{m}\frac{\delta_{j,c}}{m}-\frac{\hat{x}_{i,c}}{m}\sum_{j}^{m}\delta_{j,c}\cdot\hat{x}_{j,c}\right],$
@@
$\frac{\partial\Theta}{\partial x}$ $=\frac{\alpha}{2\left(1+\left(\frac{\pi}{2}\alpha x\right)^{2}\right)}\quad\text{where $\alpha=2$}$
@@
Figure 4: Architecture of SEW residual block proposed by fang2021deep
To quantify these sources, we analyze the gradient expressions of each component within the spike-based element-wise residual blocks. As seen in Equation 6, convolution Conv, batch normalization BN, and surrogate gradient components $\Theta$ are all dependent on input and intermediate values. However, max pooling MaxPool introduces a structural bottleneck where gradients are non-zero only at the maximum activation locations, and zero elsewhere. These null gradients can propagate through backward pass, create sparsity in earlier layers and is denoted as architectural sparsity. This effect is amplified in SEW-ResNet due to the high coupling nature of residual blocks (illustrated in Figure 4). Thus, replacing max pooling with a less sparse alternative could enhance gradient density. To evaluate this hypothesis, we replaced max pooling MaxPool operations with average pooling AvgPool, whose gradient is given by:
$\frac{\partial\textbf{AvgPool}}{\partial x_{i,j}}=\sum_{(p,q)\;:\;(i,j)\in\mathcal{R}_{p,q}}\;\frac{1}{|\mathcal{R}_{p,q}|}\cdot\frac{\partial\mathcal{L}}{\partial y_{p,q}}$ (7)
@@
Table 3: Comparison of CIFAR-100 and CIFAR-10 validation accuracy of undefended SEW-ResNet variants (fang2021deep) with average pooling. Best performance per attack on each dataset are denoted in bold. We employed similar attack settings as Table 1, accuracy reduction with respect to Table 2 is denoted in red and accuracy increment is denoted in green. There is a clear trade off in clean accuracy and adversarial robustness.
@@
CIFAR10-DVS SEW-ResNet18 46.27 ($\downarrow$0.92) 13.94 ($\downarrow$7.53) 77.77 ($\uparrow$2.59)
SEW-ResNet34 47.66 ($\downarrow$1.37) 12.54 ($\downarrow$6.11) 76.43 ($\uparrow$1.14)
SEW-ResNet50 47.01 ($\downarrow$1.51) 14.30 ($\downarrow$8.04) 78.35 ($\uparrow$1.82)
@@
As shown in Figure 5, input gradient sparsity is significantly reduced when average pooling is used in place of max pooling across all layers. Specifically, sparsity in the pooling layer is completely eliminated, leading to a 37.88% reduction compared to max pooling, while a spiking neuron layers exhibit sparsity as low as 16.52%. This observation is consistent with prior studies. For instance, fang2021deep highlighted the importance of gradient magnitude in optimizing SNNs to enhance performance, while luu2024improvement addressed training challenges by introducing signal concatenation, effectively scaling the gradient norm to improve generalization. Complete elimination of sparsity is also theoretically unachievable due to the inherent sparse nature of spiking neurons in SNNs, typically modeled using the Heaviside function (as in Equation 4) and is referred to as natural sparsity within this context.
4.3.2 Theoretical explanation
@@
Previous studies have demonstrated a strong relationship between the magnitude of weight gradients and model performance (luu2024improvement; fang2021deep; shalev2017failures). More specifically, luu2024improvement had proven the correlation between gradient scale and performance of a surrogate SNN model over gradient signal-to-noise-ratio (SNR) when scaling over temporal dimension $T$ as follow:
Theorem 4.2 (luu2024improvement).
Assume that temporally-extended gradient $g^{extend}_{\mathbf{w}}(\mathbf{x})$ does not explode nor vanish, bounded as:
$G(\mathbf{w})^{2}>\left\|\mathbb{E}_{\mathbf{x}}\left[g^{extend}_{\mathbf{w}}(\mathbf{x})\right]\right\|_{\infty}>0$
for some scalar $G(\mathbf{w})$, where both gradient of baseline SNN $g_{\mathbf{w}}(\mathbf{x})$ and temporally extended SNN $g^{extend}_{\mathbf{w}}(\mathbf{x})$ converge toward a reasonable approximation of $h(\mathbf{x})$ and have similar distribution that satisfy:
$\mathrm{Cov}\left(h(\mathbf{x}),g^{extend}_{\mathbf{w}}(\mathbf{x})\right)=0,\quad\mathrm{sign}(\mathbb{E}_{\mathbf{x}}\left[h(\mathbf{x})\right])>0,\quad h(\mathbf{x})\cdot g^{extend}_{\mathbf{w}}(\mathbf{x})\sim h(\mathbf{x})\cdot g_{\mathbf{w}}(\mathbf{x}),$
there exist:
$\mathrm{sup}\left(SNR\left(\delta^{extend}(\mathbf{x})\right)\right)\geq\mathrm{sup}\left(SNR\left(\delta(\mathbf{x})\right)\right).$
where:
$\mathrm{sup}\left(SNR(\delta^{extend}(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta^{extend}(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta^{extend}(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta^{extend}(\mathbf{x})\right]\right\|^{2}_{\infty}},$
$\mathrm{sup}\left(SNR(\delta(\mathbf{x}))\right)=\frac{\left\|\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}{\mathbb{E}_{\mathbf{x}}\left\|\delta(\mathbf{x})-\mathbb{E}_{\mathbf{x}}\left[\delta(\mathbf{x})\right]\right\|^{2}_{\infty}}$
Proof for this theorem can be found at Appendix B.
Building on this foundation, we aim to investigate how sparsity in the input gradients may suppress weight gradients during optimization, thereby influencing overall performance. To prove this statement, we consider the following theorem:
Theorem 4.3.
@@
As obtained from the derivation results, input gradient $\nabla_{s[t]}p_{\mathbf{w}}(s)$ shared a gradient chain $\frac{\partial p_{\mathbf{w}}}{\partial y_{\mathbf{w}_{2}}}\cdot\frac{\partial y_{\mathbf{w}_{2}}}{\partial z_{\mathbf{w}_{2}}}\cdot\frac{\partial z_{\mathbf{w}_{2}}}{\partial y_{\mathbf{w}_{1}}}\cdot\frac{\partial y_{\mathbf{w}_{1}}}{\partial z_{\mathbf{w}_{1}}}$ with the Jacobian $\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)$ of $\mathbf{w}_{1}$ during backward computation and a shorter chain of $\frac{\partial p_{\mathbf{w}}}{\partial y_{\mathbf{w}_{2}}}\cdot\frac{\partial y_{\mathbf{w}_{2}}}{\partial z_{\mathbf{w}_{2}}}$ with respect to $\mathbf{w}_{2}$, leading us to the initial observations. [^1] A more illustrative demonstration can be found at Figure 6
We are confident that any sparsity in the shared gradient path propagates backward to both the weight and input gradients. Still, opposite of sparsity, how can we formally quantify the relationship between the density of the weight gradient (specifically of $\mathbf{w}_{1}$, as it is being the closest parameter to input and is highly affected by the shared gradient trail) with respect to input gradient, particularly in the context of Theorem 4.3? To address this, we establish the following theorem, which provides a precise characterization of their relationship:
Theorem 4.4.
Given the same condition as Theorem 4.3, the parameter matrix $\mathbf{w}_{l}\in\mathbb{R}^{d_{1}\times d_{2}}$ at layer $l$ is said to have a row/column density degree of $k^{(i)}_{\mathbf{w}_{l}}$$k^{(j)}_{\mathbf{w}_{l}}$ (or $k^{(i)}_{\mathbf{w}_{l}}$$k^{(j)}_{\mathbf{w}_{l}}-$dense row/column for short) if each of its row/column vector $\mathbf{w}_{l}^{(i)}/\mathbf{w}_{l}^{(j)}$ (where $i\in d_{1}$, $j\in d_{2}$) has at most $k$ non-zero entries $\mathbf{w}^{(ij)}_{l}$, defined as:
@@
and all weights have $k^{(i)}_{\mathbf{w}_{l}}>0,k^{(j)}_{\mathbf{w}_{l}}>0$ post optimization, then $L_{0}$ norm of weight gradient of $\mathbf{w}_{1}$ can bounded as:
@@
$\mathbb{E}\left[\|\nabla_{s[t]}p_{\mathbf{w}}(s)\|_{0}\right]\leq k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{\mathbf{w}_{1}}$. (17)
@@
$Pr\left(\mathbf{w}^{(in)}_{2}\cdot S^{(nj)}\neq 0\right)\leq\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{d^{2}},\quad\Leftrightarrow\quad\mathbb{E}\left[\|\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)\|_{0}\right]\leq T\cdot\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{d}.$ (18)
Divide obtained inequalities from Equation 17 and 18:
$\frac{\mathbb{E}\left[\|\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)\|_{0}\right]}{\mathbb{E}\left[\|\nabla_{s[t]}p_{\mathbf{w}}(s)\|_{0}\right]}\leq\frac{T\cdot\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{d}}{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{\mathbf{w}_{1}}}$, (19)
$\Leftrightarrow$ $\mathbb{E}\left[\|\nabla_{\mathbf{w}_{1}}p_{\mathbf{w}}(s)\|_{0}\right]\leq\frac{T}{d}\cdot\frac{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{S}}{k^{(i)}_{\mathbf{w}_{2}}k^{(j)}_{\mathbf{w}_{1}}}\cdot\mathbb{E}\left[\|\nabla_{s[t]}p_{\mathbf{w}}(s)\|_{0}\right],$
@@
Combining this with observation from liu2024enhancing, we can see that there is a clear and chained connection between weight gradient, input gradient and robustness. Any attempt to regularize $L_{0}$ norm of input gradient to improve robustness as in liu2024enhancing will also dampen magnitude of weight gradient.